OpenStudy (anonymous):

Lenny’s favorite radio station has this schedule: news: 19 min. commercials: 4 min. music: 37 min. If Lenny chooses a time of day at random to turn on the radio to his favorite station, what is the probability that the news will be on?

OpenStudy (anonymous):

19/(19+4+37)=19/60

OpenStudy (anonymous):

0.31

OpenStudy (mathstudent55):

Add the three amounts: 19 min + 4 min + 37 min = 60 min Since news is 19 minutes out of every 60 minutes, the probability is \(P(news) = \frac{19}{60}\)

OpenStudy (anonymous):

thanks guys!

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

;) give a medal and become a fan